OpenStudy (anonymous):

Need quick help, Will give medal!: which polynomial identity will prove that 49 = (2+5)^2? Is it difference of squares, Difference of cubes, Sum of cubes or square of binomial?

OpenStudy (anonymous):

sqaure of binomial. Difference of anything involves subtraction, and 2/5 are both not perfect cubes

OpenStudy (solomonzelman):

\[49=7^2\]

OpenStudy (ybarrap):

(2+5)^2 = 2^2*5^0+2^1*5^1+2^0*5^2=4+10+25=49, using the binomial theorem: http://en.wikipedia.org/wiki/Binomial_theorem A binomial means two terms. The result of which is squared.